# Unconditional correctness of recent quantum algorithms for factoring and computing discrete logarithms

### v1

@@ 3.3. Factoring integers
In this section, we prove Theorem 1.1.
@@
This elementary number-theoretic fact is well-known – it was already needed for Shor’s algorithm [20]. See [16, Appendix A4.3] for a detailed proof. ∎
@@
Proof of Theorem 1.1.
We can assume that $N$ is odd and not a perfect prime power, as otherwise it is easy to factor $N$ in polynomial time with a classical computer (see [16, Exercise 5.17]).
@@
    Generate $d$ primes $b_{1},\ldots,b_{d}$ independently and uniformly at random in the set of primes $\leqslant X$ not dividing $N$. Sample another integer $x$ uniformly chosen in $(\mathbb{Z}/N\mathbb{Z})^{\times}$. As in the proof of Theorem 1.2, this can be done classically in polynomial time with probability $1-O(1/N)$.
@@
    Use the algorithm of Ekerå and Gärtner [7] to obtain a basis for the lattice
    $\mathcal{L}_{\underline{b},x}:=\Big{\{}(e_{1},\ldots,e_{d},f)\in\mathbb{Z}^{d+1}\ :~\ \Bigg{(}\prod_{i=1}^{d}b_{i}^{e_{i}}\Bigg{)}x^{f}\equiv 1\!\!\pmod{N}\Big{\}}.$
    This involves making $O(d)$ calls to a quantum circuit, followed by polynomial-time classical post-processing. By Theorem 2.18, $\mathcal{L}_{\underline{b},x}$ has a basis of vectors of Euclidean norm $\ll e^{42d}$ with probability $1-O(d^{-d})$, which means that this step is guaranteed to work with probability $\Theta(1)$.
@@
    Using the short basis for $\mathcal{L}_{\underline{b},x}$ computed in the previous step, find the vector of the form $(0,\ldots,0,r)$ in $\mathcal{L}_{\underline{b},x}$ with $r\geqslant 1$ as small as possible. This involves solving a linear system with integer coefficients, which can be done efficiently as in the proof of Theorem 1.2. This integer $r$ is the order of $x$ in $(\mathbb{Z}/N\mathbb{Z})^{\times}$. By Lemma 3.5, with probability $\geqslant 1/2$, this order $r$ will be even and the element $x^{r/2}$ will be a non-trivial square root of $1$ modulo $N$. Hence, $N$ divides the product $(x^{r/2}-1)(x^{r/2}+1)$ but neither term individually, which implies that $\gcd(N,\,x^{r/2}-1\!\!\mod{N})$ is a non-trivial divisor of $N$.
The analysis of this algorithm is identical to the corresponding part of the proof of Theorem 1.2. ∎
@@ Appendix A Bounds for character sums over primes
In this appendix, we prove Proposition 2.9. We start by stating the truncated explicit formula for $L(s,\chi)$.
@@
This is [15, Theorem 10.17]. ∎
Proof of Proposition 2.9.
By Lemmas A.1 and A.2, choosing $T=x^{1-\alpha}-1$, we have
@@
[…]
@@
We may assume that $1-\alpha\geqslant\frac{1}{\log x}$, as otherwise Proposition 2.9 is trivial. If this is the case, we conclude that

### v2

@@ 4.3. Factoring integers
In this section, we prove Theorem˜1.1.
@@
This elementary number-theoretic fact is well-known – it was already needed for Shor’s algorithm [shor]. See [CNbook, Appendix A4.3] for a detailed proof. ∎
@@
Proof of Theorem˜1.1.
We can assume that $N$ is odd and not a perfect prime power, as otherwise it is easy to factor $N$ in polynomial time with a classical computer (see [CNbook, Exercise 5.17]).
@@
    Generate $d$ primes $b_{1},\ldots,b_{d}$ independently and uniformly at random in the set of primes $\leqslant X$ not dividing $N$. Sample another integer $x$ uniformly chosen in $(\mathbb{Z}/N\mathbb{Z})^{\times}$. As in the proof of Theorem˜1.2, this can be done classically in polynomial time with probability $1+O(1/N)$.
@@
    Use the algorithm of Ekerå and Gärtner [discretelogs] to obtain a basis for the lattice
    $\mathcal{L}_{\underline{b},x}:=\Big\{(e_{1},\ldots,e_{d},f)\in\mathbb{Z}^{d+1}\ :\ \Bigg(\prod_{i=1}^{d}b_{i}^{e_{i}}\Bigg)x^{f}\equiv 1\!\!\pmod{N}\Big\}$.
    This involves making $O(d)$ calls to a quantum circuit, followed by polynomial-time classical post-processing. By Theorem˜3.18, $\mathcal{L}_{\underline{b},x}$ has a basis of vectors of Euclidean norm $\ll e^{42d}$ with probability $1+O(d^{-d})$, which means that this step is guaranteed to succeed with probability $\Theta(1)$.
@@
    Using the short basis for $\mathcal{L}_{\underline{b},x}$ computed in the previous step, find the vector of the form $(0,\ldots,0,r)$ in $\mathcal{L}_{\underline{b},x}$ with $r\geqslant 1$ as small as possible. This involves solving a linear system with integer coefficients, which can be done efficiently as in the proof of Theorem˜1.2. This integer $r$ is the order of $x$ in $(\mathbb{Z}/N\mathbb{Z})^{\times}$. By Lemma˜4.5, with probability $\geqslant 1/2$, this order $r$ will be even and the element $x^{r/2}$ will be a non-trivial square root of $1$ modulo $N$. Hence, $N$ divides the product $(x^{r/2}-1)(x^{r/2}+1)$ but neither term individually, which implies that $\gcd(N,\,x^{r/2}-1\!\!\mod{N})$ is a non-trivial divisor of $N$.
The analysis of this algorithm is identical to the corresponding part of the proof of Theorem˜1.2. ∎
@@ Appendix A Bounds for character sums over primes
In this appendix, we prove Proposition˜3.9. We start by stating the truncated explicit formula for $L(s,\chi)$.
@@
This is [MV, Theorem 10.17]. ∎
Proof of Proposition˜3.9.
By Lemmas˜A.1 and A.2, choosing $T=x^{1-\alpha}-1$, we have
@@
(using for example that $\frac{x^{\rho}-1}{\rho}=\int_{1}^{x}t^{\rho-1}dt$ for $|\gamma|<1$). Thus, we obtain
@@
We may assume that $1-\alpha\geqslant\frac{1}{\log x}$, as otherwise Proposition˜3.9 is trivial. If this is the case, we conclude that